OpenStudy (anonymous):

Jason left home at 0830 and cycled at an average speed of 3m/s from his house to the bookshop. On reaching the bookshop at 0840, he browsed and bought two books 25 minutes later. He doubled his speed on the return journey. What is his average speed for the entire journey in km/h?

OpenStudy (anonymous):

Time taken = 10 minutes Speed = 3m/s Distance = s * t = 600 * 3 = 1800 m Total distance = 1800 * 2 = 3600 m Total time = 10 + 5 = 15 min. = 15 * 60 = 900 s Average speed = 3600 m / 900 s = 4 m/s

OpenStudy (anonymous):

Er.. What is 4m/s in km/h?

OpenStudy (anonymous):

4 * 18/5 = 14.4 km/hr.

OpenStudy (anonymous):

thx a lot